OpenStudy (anonymous):

Megan is four years younger than her aunt, and the sum of their ages is at least 40. Which inequality would you use to solve determine the ages of Megan and her aunt? * a - 4 + a ≤ 40 a - 4 = 40 2a - 4 ≥ 40 m + a - 4 > 40

OpenStudy (anonymous):

a - 4 + a ≤ 40

OpenStudy (dmndlife24):

@gerryliyana is right It's the only inequality that 40 is greater than or equal too.

OpenStudy (anonymous):

thanx can u help wit some more

OpenStudy (anonymous):

Megan is four years younger than her aunt is equal to M = A - 4 so the sum of their ages : M + A = A - 4 + A

OpenStudy (anonymous):

you are welcome :)

OpenStudy (anonymous):

no,2a-40 >=40 is correct.

OpenStudy (anonymous):

sum of their is at least 40 means sum of their ages is 40 or more.

OpenStudy (anonymous):

Translate "the product of negative four and the sum of ten and a number" into a variable expression. * -4(10 + n) -4 / 10 + n -4(10n) -4 + (10 + n)

OpenStudy (anonymous):

sum of 10 and n=10+n find product with -4

OpenStudy (anonymous):

a

OpenStudy (anonymous):

like what i said